# Language Model Agents Suffer from Compositional Generalization in Web Automation

### v1

@@ Abstract
Language model agents (LMA) recently emerged as a promising paradigm on muti-step decision making tasks, often outperforming humans and other reinforcement learning agents. Despite the promise, their performance on real-world applications that often involve combinations of tasks is still underexplored. In this work, we introduce a new benchmark, called CompWoB – 50 new compositional web automation tasks reflecting more realistic assumptions. We show that while existing prompted LMAs (gpt-3.5-turbo or gpt-4) achieve 94.0% average success rate on base tasks, their performance degrades to 24.9% success rate on compositional tasks. On the other hand, transferred LMAs (finetuned only on base tasks) show less generalization gap, dropping from 85.4% to 54.8%. By balancing data distribution across tasks, we train a new model, HTML-T5++, that surpasses human-level performance (95.2%) on MiniWoB, and achieves the best zero-shot performance on CompWoB (61.5%). While these highlight the promise of small-scale finetuned and transferred models for compositional generalization, their performance further degrades under different instruction compositions changing combinational order. In contrast to the recent remarkable success of LMA, our benchmark and detailed analysis emphasize the necessity of building LMAs that are robust and generalizable to task compositionality for real-world deployment.
@@ 1 Introduction
Based on the exceptional capability of large language models (LLMs) (OpenAI, 2023; Anil et al., 2023; Touvron et al., 2023) in commonsense understanding (Brown et al., 2020; Chowdhery et al., 2022), multi-step reasoning (Wei et al., 2022; Kojima et al., 2022), program synthesis (Chen et al., 2021) and self-improvement (Shinn et al., 2023; Madaan et al., 2023; To et al., 2023), language model agents (LMA) have recently emerged to tackle various decision making problems, such as robotics (Huang et al., 2022a; Ahn et al., 2022), information retrieval (Nakano et al., 2021; Yao et al., 2022b), and external tool use (Wu et al., 2023; Shen et al., 2023; Lu et al., 2023). Especially, in web automation (Shi et al., 2017), LMAs with prompting (Kim et al., 2023; Sun et al., 2023; Zheng et al., 2023) outperform humans and other learning-based agents, such as reinforcement learning (Humphreys et al., 2022) or finetuned language models (Gur et al., 2022; Furuta et al., 2023).
@@
Despite their proficiency in MiniWoB (Shi et al., 2017), a standard web automation benchmark, it is still unclear whether LMAs could deal with challenges in the real world: such as complex observation (Gur et al., 2023), domain generalization (Deng et al., 2023), and ambiguity of instructions (Zhou et al., 2023b). These challenges are exacerbated due to the open-ended nature of real-world tasks, making it infeasible to prepare exemplars and prompts in advance for any unseen task.
@@
In this work, we extensively study the generalization of LMAs to more realistic task compositions. We first design a new controlled test bed, called CompWoB, with 50 compositional tasks by combining a set of base tasks based on their difficulty (Figure 1). Each compositional task is implemented from 2 to 8 base tasks in a single-page or multi-page environment with instructions linked together using simple connectors such as “and then”. Only providing the knowledge about base tasks, we investigate the generalization performance of existing SoTA prompted LMAs (Kim et al., 2023; Sun et al., 2023; Zheng et al., 2023) with planning, self-improvement, program synthesis, and structured prompts that are supported by gpt-3.5-turbo and gpt-4. Our findings indicate that their performance drops significantly, from 94.0% success on base tasks to 24.9% success on compositional tasks. In contrast, small-scale LMAs finetuned only on base tasks and zero-shot-transferred to compositional settings (i.e. transferred LMAs), deal with unknown task compositionality better, achieving 54.8% success rate on average. By rebalancing the data distribution, we also train a new model, HTML-T5++, that achieves human-level performance on MiniWob and performs the best among all the LMAs on compositional tasks. We further show that LMAs struggle to handle complex instruction compositions permuting the order of sub-instructions, where prompted agents are more robust to the difference in the order of compositions compared to transferred agents (6.9% vs 23.8% drop in performance). Finally, we illustrate that instruction length and observation complexity are useful indicators of compositional task performance.
@@
Figure 1: We design CompWoB, as a novel decision making benchmark for LMAs, by leveraging the high composability of simulated web environments. We first select base tasks from the original MiniWoB (Shi et al., 2017) based on the brute-force task complexity averaged among existing LMAs. While considering the feasibility, we randomly combine them into a single task (e.g. click-checkboxes-transfer + enter-password + click-dialognormal-$\rightarrow$ click-checkboxes-transfer_enter-password_click-dialog). The instructions of base tasks are stitched with “and then”. LMAs are asked to satisfy the given instructions sequentially (e.g. satisfying the success criteria of task A $\rightarrow$ B $\rightarrow$ C). We also implement reverse-order instruction settings, where the instructions are provided upside down (e.g. solve task B, and solve task C, after solving task A). These complex yet controllable strategies make the analysis of LMA’s behaviors easy, while maintaining complex and ambiguous aspects of real-world web environments.
@@
  • We develop CompWoB  [^1] https://github.com/google-research/google-research/tree/master/compositional_rl/compwob , simulated web environments for LMAs to measure the generalization to the realistic task compositionality and complex instructions.
  • We propose a new data mixture strategy for finetuning LMA. HTML-T5++, trained on this strategy, achieves human-level performance on MiniWoB (95.2%) and the best zero-shot transfer to CompWoB (61.5%).
@@ 2 Related Works
Language Model Agents  Beyond the common NLP tasks, LLMs could act as autonomous agents (Wang et al., 2023b; Qin et al., 2023a) to solve the given instruction-following tasks, by considering the context in the prompt as states (Ahn et al., 2022; Yao et al., 2022b; Hsieh et al., 2023) and sequentially planning and manipulating external “tools” or “actuators”, such as calculators (Parisi et al., 2022), retrievers (Schick et al., 2023; Hao et al., 2023), APIs (Qin et al., 2023b; Tang et al., 2023a), programs (Gao et al., 2023; Wang et al., 2023a; Liang et al., 2023; Song et al., 2023; Cai et al., 2023), robotic commands (Huang et al., 2022a; b; Tang et al., 2023b), computer game (Nottingham et al., 2023; Wang et al., 2023c), or other foundation models (Lu et al., 2023; Hsieh et al., 2023; Wu et al., 2023; Shen et al., 2023; Yang et al., 2023). Those prior works have worked on proposing novel benchmarks (Li et al., 2023; Xu et al., 2023; Patil et al., 2023) and comparing backbone LLMs (e.g. open-sourced v.s. private) (Ruan et al., 2023; Liu et al., 2023b; a). Despite their success, it is still unclear how such LMAs designed for specific tasks can generalize out-of-domain problems, which should be an important perspective since we may not prepare prompts and exemplars for all the possible combinations of problems in the real world. In this work, we measure the compositional generalization and robustness to the complex instructions in web automation.
@@
Web Automation  Although prior works have worked on imitation learning and reinforcement learning (Liu et al., 2018; Gur et al., 2019; Jia et al., 2019; Humphreys et al., 2022), web automation has become a popular domain as a benchmark for LMAs (Gur et al., 2022; Kim et al., 2023). In earlier work, finetuned LMAs, based on at most 3-billion parameters, amortize the training costs with the strong prior knowledge on web environments (Gur et al., 2022; Furuta et al., 2023; Shaw et al., 2023), but they often result in sub-optimal performances due to the insufficient data coverage. Recently, by leveraging capable private LLMs (Brown et al., 2020; Ouyang et al., 2022) with self-refinement (Kim et al., 2023), program synthesis (Sun et al., 2023), well-designed structured prompts with instruction-state translation (Zheng et al., 2023), or hierarchical prompts (Sridhar et al., 2023; Ma et al., 2023), prompted LMAs with few-shot exemplars have outperformed finetuned LMAs and shown competitive performance to humans and RL-finetuned agents. In contrast, our work discusses the generalization and robustness of those LMAs in zero-shot web automation with a set of compositional tasks, and resolves the sub-optimality of finetuned LMAs via data-rebalancing.
@@
In addition to MiniWoB (Shi et al., 2017), a representative web simulator, several works have conducted real-world evaluation (Gur et al., 2023) and proposed novel benchmarks reflecting real-world assumptions, such as a simulated e-commerce site (Yao et al., 2022a), sand-boxed real-world websites (Zhou et al., 2023b), an adaptive sim-to-real bridge with unsupervised auto-curricula (Gur et al., 2021), and large-scale web interaction dataset curated by human annotators (Deng et al., 2023). However, real-world web automation may make the analysis challenging because it often faces many obstacles, such as complex HTML observations, domain gaps between websites, and ambiguous instructions. In this work, we design CompWoB while controlling task difficulty and ambiguity of instructions, and investigate what factors may prevent the generalization capability of LMAs in compositional tasks.
@@
Large Language Models for Compositional Tasks  Several works have investigated compositional natural language problems with LLMs, such as semantic parsing (Furrer et al., 2021; Shaw et al., 2021; Zhou et al., 2023a), logic grid puzzles (Dziri et al., 2023), mathematical reasoning (Chen et al., 2023), programming (Zelikman et al., 2023), and planning (Brahman et al., 2023), which shows that dynamical selection of exemplars for decomposed sub-problems (Drozdov et al., 2023) or model scaling (Qiu et al., 2022) could help generalization. While those are focused on static tasks, our paper studies compositional generalization in decision making, especially, in web automation where the task may have more explicitly decomposable structures (Gur et al., 2021) than natural language tasks.
@@ 3 Preliminaries
Web automation could be described as a deterministic sequential decision making problem, which consists of a state space $\mathcal{S}$, action space $\mathcal{A}$, deterministic transition function $T:\mathcal{S}\times\mathcal{A}\xrightarrow{}\mathcal{S}$, a set of instructions $\mathcal{G}$, a set of contextual information (i.e. prompts for LLM) $\mathcal{C}$, and episodic reward function (i.e. success criteria) $r:\mathcal{S}\times\mathcal{G}\times\mathcal{A}\xrightarrow{}\{0,1\}$. At each time step $t$, the language model agent $\pi$ infers the action conditioned on the prompt, instruction, current state, and previous actions $\pi:\mathcal{S}\times\underbrace{\mathcal{A}\times\cdots\times\mathcal{A}}_{\times t}\times\mathcal{C}\times\mathcal{G}\rightarrow{}\mathcal{A}$, and moves to the next state: $s_{t+1}=T(s_{t},a_{t})$. When the agent reaches the terminal state (e.g. Login button is clicked) or the max time step is exceeded, the episode is marked as a success if the instruction $g$ is satisfied (i.e. $r(s_{t},g,a_{t})=1$). The state $s_{t}\in\mathcal{S}$ is a raw HTML, and we assume the programmatic action space: function(selector, text). function is either click, move or type, selector is an integer index or XPath that can uniquely specify the element, and text is a text input for type function.
Task Compositionality  Web automation tasks can be decomposed into a set of primitive base tasks. For instance, (1) clicking several checkboxes, (2) fulfilling the password form, and (3) closing the dialog window. Such a combination could be open-ended. In this work, we assume that the task $\psi\in\Psi$ is characterized by a corresponding subtree of HTML ($\mathcal{S}_{\psi}\subset\mathcal{S}$) and instructions ($\mathcal{G}_{\psi}\subset\mathcal{G}$), and can be combined each other as long as the task is feasible and executable.
@@ 4 Language Model Agents for Web Automation
We here review the existing LMAs for web automation, such as RCI (Section 4.1), AdaPlanner (Section 4.2), Synapse (Section 4.3), and transferred LMAs (Section 4.4). To clarify their algorithmic difference, we provide the pseudo code for prompted LMAs in Appendix B. Moreover, we resolve the sub-optimal performance of transferred LMAs with data-rebalanced finetuning (Section 4.5).
@@ 4.1 RCI
The agent with Recursive Criticism and Improvement (RCI) prompting (Kim et al., 2023) first generates an open-loop plan to follow a given instruction using few-shot demonstrations. Next, it uses a prompt-based critic to identify the problems in the plan and improves the plan by reflecting on self-criticized outputs, which is referred as an explicit RCI (ERCI) loop. After ERCI, the agent follows the self-improved plan step-by-step. Before executing the action at each time step, the agent grounds the action to the current state (i.e. HTML, open-loop plan, and previous actions) and refines its formatting to be parsable, which increases the feasibility and reduces hallucinations. These final steps are referred to as an implicit RCI (IRCI) loop without the self-criticism.
All of those play complementary roles to achieve proficient performance. While 1 ERCI and 3 IRCI loops are recommended by Kim et al. (2023), we observe that the optimal number of self-improvement iterations may differ across the tasks. See Appendix C.1 for the details.

### v3

# Exposing Limitations of Language Model Agents in Sequential-Task Compositions on the Web

@@ Abstract
Language model agents (LMA) recently emerged as a promising paradigm on muti-step decision making tasks, often outperforming humans and other reinforcement learning agents. Despite the promise, their performance on real-world applications that often involve combinations of tasks is still underexplored. In this work, we introduce a new benchmark, called CompWoB – 50 new compositional web automation tasks reflecting more realistic assumptions. We show that while existing prompted LMAs (gpt-3.5-turbo or gpt-4) achieve 94.0% average success rate on base tasks, their performance degrades to 24.9% success rate on compositional tasks. On the other hand, transferred LMAs (finetuned only on base tasks) show less generalization gap, dropping from 85.4% to 54.8%. By balancing data distribution across tasks, we train a new model, HTML-T5++, that surpasses human-level performance (95.2%) on MiniWoB, and achieves the best zero-shot performance on CompWoB (61.5%). While these highlight the promise of small-scale finetuned and transferred models for task compositionality, their performance further degrades under different instruction compositions changing combinational order. In contrast to the recent remarkable success of LMA, our benchmark and detailed analysis emphasize the necessity of building LMAs that are robust and generalizable to task compositionality for real-world deployment.
@@ 1 Introduction
Based on the exceptional capability of large language models (LLMs) (OpenAI, 2023; Anil et al., 2023; Touvron et al., 2023) in commonsense understanding (Brown et al., 2020; Chowdhery et al., 2022), multi-step reasoning (Wei et al., 2022; Kojima et al., 2022), program synthesis (Chen et al., 2021) and self-improvement (Shinn et al., 2023; Madaan et al., 2023; To et al., 2023), language model agents (LMAs) have recently emerged to tackle various decision making problems, such as robotics (Huang et al., 2022a; Ahn et al., 2022), information retrieval (Nakano et al., 2021; Yao et al., 2022b), and external tool use (Wu et al., 2023; Shen et al., 2023). Especially, web automation (Shi et al., 2017) has attracted attention a lot as an promising application of LMAs, because LMAs with prompting (Kim et al., 2023; Sun et al., 2023; Zheng et al., 2023) outperform humans and other learning-based agents, such as imitation (Furuta et al., 2023) or reinforcement learning (Humphreys et al., 2022).
@@
Despite their human-level proficiency in MiniWoB (Shi et al., 2017), a standard web automation benchmark, it is still unclear whether LMAs could deal with challenges in the real world; such as complex observation (Gur et al., 2023), domain generalization (Deng et al., 2023), and ambiguity of instructions (Zhou et al., 2023b). These challenges are exacerbated due to the open-ended nature of real-world tasks, making it infeasible for the agent system to prepare exemplars and prompts for any unseen task in advance. Moreover, it is reasonable and practically important to assume the zero-shot evaluation of the off-the-shelf agents at deployment. This is an orthogonal assumption to the recent progress in LMAs learnable from execution/self-feedback (Shinn et al., 2023; Ma et al., 2023) during the deployment. Considering the agents are deployed as a service, (1) it is infeasible to cover all the possible prompts/demonstrations for user requests in advance and (2) iterative trial-and-error to adjust system prompts hurts user experiences.
@@
In this work, we extensively study the transferability of LMAs to more realistic sequential task compositions. We first design a new controlled test bed, called CompWoB, with 50 compositional tasks by combining a set of base tasks based on their difficulty (Figure 1, right). CompWoB works as a hold-out test environment accompanying with MiniWoB as a train environment (Figure 1, left). Each compositional task is implemented from 2 to 8 base tasks in a single-page or multi-page environment with instructions linked together using simple connectors such as “and then”. Only providing the knowledge about base tasks, we investigate the generalization performance of existing SoTA prompted LMAs (Kim et al., 2023; Sun et al., 2023; Zheng et al., 2023) with planning, self-improvement, program synthesis, and structured prompts that are supported by gpt-3.5-turbo and gpt-4. Our findings indicate that their performance drops significantly, from 94.0% success on base tasks to 24.9% success on compositional tasks. In contrast, small-scale LMAs finetuned only on base tasks and zero-shot-transferred to compositional settings (i.e. transferred LMAs), deal with unknown task compositionality better, achieving 54.8% success rate on average. By rebalancing the data distribution, we train a new model, HTML-T5++, that achieves human-level performance on MiniWoB and performs the best among all the LMAs on compositional tasks. In contrast, small-scale LMAs finetuned only on base tasks and zero-shot-transferred to compositional settings (i.e. transferred LMAs), deal with unknown task compositionality better, achieving 54.8% success rate on average. By rebalancing the data distribution, we also train a new model, HTML-T5++, that achieves human-level performance on MiniWoB and performs the best among all the LMAs on compositional tasks. We further point out that LMAs struggle to handle complex instruction compositions permuting the order of sub-instructions, where prompted agents are more robust to the difference in the order of compositions compared to transferred agents (6.9% vs 23.8% drop in performance). Finally, we illustrate that instruction length and observation complexity are useful indicators of compositional task performance.
@@
Figure 1: (Left) We design CompWoB, as a novel decision making benchmark for LMAs, by leveraging the high composability of simulated web environments. We first select base tasks from the original MiniWoB (Shi et al., 2017) based on the brute-force task complexity averaged among existing LMAs. While considering the feasibility and reality, we systematically combine them into a sequential single task (e.g. click-checkboxes-transfer + enter-password + $\rightarrow$ click-checkboxes-transfer_enter-password_click-dialog). The instructions of base tasks are stitched with “and then”. LMAs are asked to satisfy the given instructions sequentially (e.g. satisfying the success criteria of task A $\rightarrow$ B $\rightarrow$ C). We also implement reverse-order instruction settings, where the instructions are provided upside down (e.g. solve task B, and solve task C, after solving task A). These complex yet controllable strategies make the analysis of LMA’s behaviors easy, while maintaining complex and ambiguous aspects of real-world web environments. (Right) In this paper, we assume the zero-shot evaluation of the off-the-shelf agents at deployment, which is practically important because (1) it is infeasible to cover all the possible prompts/demonstrations for user requests in advance and (2) iterative trial-and-error to adjust system prompts hurts user experiences. CompWoB works as a hold-out test environment accompanying with MiniWoB as a train environment.
@@
  • We develop CompWoB  [^1] https://github.com/google-research/google-research/tree/master/compositional_rl , simulated web environments for LMAs to measure the generalization to the realistic task compositionality and complex instructions.
  • We explore a new data mixture heuristic for finetuning LMAs, where we find that it is effective to add demonstrations of the tasks that suffer from data shortage and then gradually reduce the ratio of easier tasks to focus more on challenging tasks. HTML-T5++, trained on our novel heuristic, achieves human-level performance on MiniWoB (95.2%) and the best zero-shot transfer to CompWoB (61.5%).
  • Our experiments reveal that while more capable models, such as gpt-4, lead to better performance, they still struggle with task compositionality or complex instructions.
@@ 2 Related Works
Web Automation  Although prior works have worked on imitation learning and reinforcement learning (Liu et al., 2018; Gur et al., 2019; Jia et al., 2019; Humphreys et al., 2022), web automation has become a popular domain as an application of LMAs (Gur et al., 2022; Kim et al., 2023). In earlier work, finetuned LMAs, based on at most 3-billion parameters, amortize the training costs with the strong prior knowledge on web environments (Gur et al., 2022; Furuta et al., 2023; Shaw et al., 2023), but they often result in sub-optimal performances due to the insufficient data coverage. Recently, by leveraging capable private LLMs (Brown et al., 2020; Ouyang et al., 2022) with self-refinement (Kim et al., 2023), program synthesis (Sun et al., 2023), structured instruction-state translation (Zheng et al., 2023), or hierarchical prompts (Sridhar et al., 2023; Ma et al., 2023), prompted LMAs with few-shot exemplars have outperformed finetuned LMAs and shown superior performance to humans and RL-finetuned agents. In contrast, our work discusses the transferability and robustness of those LMAs in zero-shot web automation with a set of compositional tasks, and resolves the sub-optimality of finetuned LMAs via data-rebalancing.
@@
In addition to MiniWoB (Shi et al., 2017), a representative web simulator, several works have conducted real-world evaluation (Gur et al., 2023) and proposed novel benchmarks reflecting real-world assumptions, such as a simulated e-commerce site (Yao et al., 2022a), sand-boxed real-world websites (Zhou et al., 2023b; Drouin et al., 2024), an adaptive sim-to-real bridge with unsupervised auto-curricula (Gur et al., 2021), and large-scale web interaction dataset curated by human annotators (Deng et al., 2023). However, real-world web automation may make the analysis challenging because it often faces miscellaneous obstacles, such as complex HTML observations, domain gaps between websites, and ambiguous instructions. In this work, we design CompWoB under realistic assumptions while controlling task difficulty and ambiguity of instructions, and investigate what may prevent the generalization capability of LMAs in compositional tasks.
Language Model Agents   Beyond the common NLP tasks, LLMs could act as autonomous agents (Wang et al., 2023b; Qin et al., 2023a) to solve the given instruction-following tasks, by considering the context in the prompt as states (Ahn et al., 2022; Yao et al., 2022b; Hsieh et al., 2023) and sequentially planning and manipulating external “tools” or “actuators”, such as calculators (Parisi et al., 2022), retrievers (Schick et al., 2023; Hao et al., 2023), APIs (Qin et al., 2023b; Tang et al., 2023a), programs (Gao et al., 2023; Wang et al., 2023a; Liang et al., 2023; Song et al., 2023; Cai et al., 2023), robotic commands (Huang et al., 2022a; b; Tang et al., 2023b), computer game (Nottingham et al., 2023; Wang et al., 2023c), or other foundation models (Lu et al., 2023; Hsieh et al., 2023; Wu et al., 2023; Shen et al., 2023; Yang et al., 2023). Those prior works have worked on proposing novel benchmarks (Li et al., 2023; Xu et al., 2023; Patil et al., 2023) and comparing backbone LLMs (e.g. open-sourced v.s. private) (Ruan et al., 2023; Liu et al., 2023b; a). Despite their success, it is still unclear how such LMAs designed for specific tasks can generalize out-of-domain problems, which should be an important perspective since we may not prepare prompts and exemplars for all the possible problems in the real world. In this work, we measure the transferability and robustness to unseen compositionality and instructions in a realistic web automation scenario.
@@
Large Language Models for Compositional Tasks  Several works have investigated compositional natural language problems with LLMs, such as semantic parsing (Furrer et al., 2021; Shaw et al., 2021; Zhou et al., 2023a), logic grid puzzles (Dziri et al., 2023), mathematical reasoning (Chen et al., 2023), programming (Zelikman et al., 2023), and planning (Brahman et al., 2023), which shows that dynamical selection of exemplars for decomposed sub-problems (Drozdov et al., 2023) or model scaling (Qiu et al., 2022) could help generalization. While those are focused on static tasks, our paper deals with task compositionaliy in interactive decision making, especially, in web automation where the task may have more explicitly decomposable structures (Gur et al., 2021) than natural language tasks.
@@ 3 Preliminaries
Web automation could be described as a deterministic sequential decision making problem, which consists of a state space $\mathcal{S}$, action space $\mathcal{A}$, deterministic transition function $T:\mathcal{S}\times\mathcal{A}\xrightarrow{}\mathcal{S}$, a set of instructions $\mathcal{G}$, a set of contextual information (i.e. prompts for LLM) $\mathcal{C}$, and episodic reward function (i.e. success criteria) $r:\mathcal{S}\times\mathcal{G}\times\mathcal{A}\xrightarrow{}\{0,1\}$. At each time step $t$, the language model agent $\pi$ infers the action conditioned on the prompt, instruction, current state, and previous actions $\pi:\mathcal{S}\times\mathcal{A}^{\times t}\times\mathcal{C}\times\mathcal{G}\rightarrow{}\mathcal{A}$, and moves to the next state: $s_{t+1}=T(s_{t},a_{t})$. When the agent reaches the terminal state (e.g. Login button is clicked) or the max time step is exceeded, the episode is marked as a success if the instruction $g$ is satisfied (i.e. $r(s_{t},g,a_{t})=1$). The state $s_{t}\in\mathcal{S}$ is a raw HTML, and we assume the programmatic action space: function(selector, text). function is either click, move or type, selector is an integer index or XPath that can uniquely specify the element, and text is a text input for type function.
Task Compositionality  Web automation tasks can be decomposed into a set of primitive base tasks. For instance, (1) clicking several checkboxes, (2) fulfilling the password form, and (3) closing the dialog window. Such a combination could be open-ended and have some dependencies. In this work, we assume that the task $\psi\in\Psi$ is characterized by a corresponding subtree of HTML ($\mathcal{S}_{\psi}\subset\mathcal{S}$) and instructions ($\mathcal{G}_{\psi}\subset\mathcal{G}$), and can be dependently combined each other as long as the compositional task is executable.
@@ 4 Language Model Agents for Web Automation
To be self-contained, we here review the representative LMAs, which are selected based on their superior performance and novelty in using LLMs for web automation; RCI (Section 4.1), AdaPlanner (Section 4.2), Synapse (Section 4.3). These are characterized with different base LLMs, prompting and pipeline. To clarify their methodological difference, we provide the pseudo code in Appendix B. Moreover, we resolve the sub-optimal performance of transferred LMAs (Section 4.4) with data-rebalanced finetuning (Section 4.5).
@@ 4.1 RCI
The agent with Recursive Criticism and Improvement (RCI) prompting (Kim et al., 2023) first generates an open-loop plan to follow a given instruction using few-shot demonstrations. Next, it uses a prompt-based critic to identify the errors in the plan and improves it by reflecting on self-criticized outputs, which is referred as an explicit RCI (ERCI) loop. After ERCI, the agent follows the self-improved plan step-by-step. Before executing the action at each step, the agent grounds the action to the current state (i.e. HTML, open-loop plan, and previous actions) and refines its formatting to be parsable, which increases the feasibility and reduces hallucinations. These final steps are referred as an implicit RCI (IRCI) loop without the self-criticism. All of those play complementary roles to achieve human-level proficiency. While 1 ERCI and 3 IRCI loops are recommended, we observe that the optimal number of self-improvement iterations may differ across the tasks (see Appendix C.1).
4.2 AdaPlanner
In contrast to other LMAs, AdaPlanner (Sun et al., 2023) leverages the capability of program synthesis in LLMs to mitigate the hallucination in a plan. Conditioning on the instruction, the description of permissible actions in the web environments, and few-shot demonstrations, the agent first generates an open-loop plan in a Python function, where each snippet corresponds to the action. Once the agent receives environmental feedback at each step, such as assertion errors in the code, other functional errors, or “ask” action to LLMs, it adaptively re-generates the plan for the remaining steps in a closed-loop manner. LLMs more capable of code have performed better, such as text-davinci-003 than gpt-3.5-turbo.